OpenStudy (anonymous):

PLEASE HELP ME I WILL MEDAL AND FAN!!!!!!!!!!!!!!! Which rate is the best deal for a bottle of mustard? A. $1.26 for 6 oz B. $1.65 for 8 oz C. $2.42 for 12 oz D. $3.30 for 16 oz

OpenStudy (jhannybean):

Just divide each amount by the cost to get the cost of each individual one

OpenStudy (jhannybean):

\[\frac{1.26}{6}=~?\]\[\frac{1.65}{8}=~?\]\[\frac{2.42}{12}=~?\]\[\frac{3.30}{16}=~?\]

OpenStudy (littlebird):

Whichever one of those is smallest is the answer

OpenStudy (jhannybean):

So once you figure out the price of each... what @Littlebird said, lol

OpenStudy (anonymous):

tnx but does - mean division

OpenStudy (jhannybean):

- = negative = subtraction

OpenStudy (littlebird):

If you see a line with a number on top and underneath it, you divide

OpenStudy (jhannybean):

Oh,that way thats a big line. I thought you were referring to something like this: 4 - 3 = 1. \(\uparrow\) That line.

OpenStudy (anonymous):

tnx @Littlebird